# Fluctuations for non-Hermitian dynamics

### v1

@@ 4 Independence of singular values
For fixed parameters $0<\omega_{K}\ll\tilde{\omega}$, $K=N^{\omega_{K}}$ we have
$\left|\frac{{\rm d}}{{\rm d}t}\langle b_{i}^{s}-b_{i}^{r}\rangle_{t}\right|\leqslant N^{-\tilde{\omega}}\mathds{1}_{|i|,|j|\leqslant K}+4(1-\mathds{1}_{|i|,|j|\leqslant K}).$
We denote the empirical particle density of the initial data $\bm{s}$ as $\mu_{0}=\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\delta_{s_{i}(0)},$ and it Stieltjes transform by $m_{0}(z)$ [^4] Within this section we use $m(z)$ to denote the Stieljes transforms of a measure $\mu$. Instead, we point out that in Theorem 4.1 we used $s^{z}(w)$ to denote the Stieltjes transform of limiting distribution of the singular values of $X-z$.. We assume that $\mu_{0}$ is regular in the following sense.
Assumption 4.3 ($(\nu,G)$-regularity).
Fix $\nu\in(0,1],G>0$. Let $\widetilde{\mu}$ be a probability measure with Stieltjes transform $\widetilde{m}$ satisfying, for some fixed $C>0$,
$|\widetilde{m}(z)|,|\widetilde{m}^{\prime}(z)|\leqslant C,\quad C^{-1}\leqslant{\rm Im}[\widetilde{m}(z)]\leqslant C\ \text{ for all }\ |z|\leqslant G.$ (4.8)
We say that $\mu$ is $(\nu,G)$-regular if, for such a $\tilde{\mu}$, its Stieltjes transform $m$ satisfies
$|m(z)-\widetilde{m}(z)|\leqslant\frac{\varphi}{\sqrt{N{\rm Im}[z]}}\ \ \mbox{for any $|z|\leqslant G$ with ${\rm Im}[z]\geqslant\frac{\varphi}{N}$}\normalcolor.$ (4.9)
@@
For our later use, we may deal with the case that $|\widetilde{m}_{0}(z)|,|\widetilde{m}_{0}^{\prime}(z)|$ may grow with $N$ or ${\rm Im}[\widetilde{m}_{0}(z)]$ may vanishes slowly with $N$, namely
$|\widetilde{m}_{0}(z)|,|\widetilde{m}_{0}^{\prime}(z)|\leqslant N^{\operatorname{o}(1)},{\rm Im}[\widetilde{m}_{0}(z)]\geqslant N^{-\operatorname{o}(1)}.$
@@
Let the processes $\bm{s}(t),\bm{r}(t)$ be solutions of the stochastic differential equations (4.5) and (4.6). Assume that the driving continuous martingales satisfy Assumption 4.2. Additionally assume that the initial data $\bm{s}(0)=\bm{r}(0)$ satisfies $(\nu,G)$-regularity in the sense of Assumption 4.3.
Then, for any choice of small constants $\nu,\omega\ll\omega_{t}\ll\omega_{K}\ll\tilde{\omega}$ and $t=N^{-1+\omega_{t}}$, with very high probability
$|s_{i}(t_{f})-r_{i}(t_{f})|\leqslant N^{-1-\omega},\ \mbox{for all $|i|\leqslant N^{\omega}$}.$
@@
Let the processes $\bm{s}(t),\bm{s}^{\prime}(t)$ both satisfy the stochastic differential equation (4.5), with different initial data and the same driving processes $b_{i}$, which are arbitrary continuous martingales (satisfying (4.7)) and $\bm{s}(0),\bm{s}^{\prime}(0)$ are $(\nu,G)$-regular around $\widetilde{\mu}_{0},\widetilde{\mu}^{\prime}_{0}$ in the sense of Assumption 4.3. Then for any $\varepsilon>0$ and $t\geqslant N^{-1+\omega_{t}}$ with $\omega_{t}\gg\nu$, for any $|i|\leqslant N$, with very high probability one has
$|\widetilde{\rho}_{t}(0)s_{i}(t)-\widetilde{\rho}^{\prime}_{t}(0)s^{\prime}_{i}(t)|\leqslant N^{\varepsilon}\cdot\frac{|i|}{N}\cdot\left(\frac{1}{\sqrt{Nt}}+\max(\frac{|i|}{N},t)\right),$ (4.10)
where $\widetilde{\rho}_{t}$ (resp. $\widetilde{\rho}^{\prime}_{t}$) is the density of $\widetilde{\mu}_{0}\boxplus\mu_{\rm sc}^{(t)}$ (resp. $\widetilde{\mu}_{0}^{\prime}\boxplus\mu_{\rm sc}^{(t)}$) (see the discussion in Section 4.4).
@@
We first prove that (4.4) holds for any $t\in[N^{-1+\omega_{t}},2N^{-1+\omega_{t}}]$ for some $\omega,\omega_{t}>0$.
@@
${\mathrm{d}}X_{s}=\frac{{\mathrm{d}}B_{s}}{\sqrt{N}},\qquad X_{0}=Y,$ (4.11)
with $Y=(1-t)^{1/2}X$ and $X$ being an i.i.d. matrix satisfying Assumption 2.1. Fix $z_{l}$, for $l=1,2$, and denote the Hermitization of $X_{s}-z_{l}$ by $H_{s}^{z_{l}}$. Then (4.11) induces the following flow on the eigenvalues $x_{i}^{z_{l}}(s)$ of $H_{s}^{z_{l}}$:
@@
with $b_{-i}^{z_{l}}(s)=-b_{i}^{z_{l}}(s)$. Let ${\bm{w}}_{i}^{z_{l}}(s)$ be the eigenvectors of $H^{z_{l}}(s)$ (they have the form (3.4)), then, for $1\leqslant i,j\leqslant N$, we have
@@
Then, by (4.3), it follows that
$\frac{{\rm d}}{{\rm d}s}\langle b_{i}^{z_{1}},b_{j}^{z_{2}}\rangle_{s}\leqslant N^{-2\tilde{\omega}},\qquad\quad 1\leqslant i,j\leqslant N^{\omega_{K}},$
with very high probability. This shows that $\{b_{i}^{z_{1}}\}$, $\{b_{i}^{z_{2}}\}$ satisfy Assumption 4.2. Additionally, by [CipErdSch2021, Theorem 3.1] we readily see that the initial condition of (4.11) satisfies Assumption 4.3. We can thus apply Propositions 4.54.6 to see that
$\big{|}\widetilde{\rho}_{t}(0)x_{i}^{z_{l}}(t)-\rho_{\mathrm{sc}}^{(t)}(0)y_{i}^{(l)}(t)\big{|}\leqslant N^{-1-\omega},\qquad\quad|i|\leqslant N^{\omega},$
with very high probability. Here $y_{i}^{(l)}(t)$, with $l=1,2$, are two fully independent processes, with $y_{i}^{(l)}(t)$ the eigenvalues of the Hermitization of a Ginibre matrix, satisfying (4.12) with $b_{i}^{z_{l}}(t)$ being replaced by $\beta_{i}^{(l)}$, where $\beta_{i}^{(1)}$, $\beta_{i}^{(2)}$ are a family of $2N$ i.i.d. standard real Brownian motions.
The transfer from this estimate on the (signed) eigenvalues $(x_{i},y_{i})$ to the singular values $(\lambda_{i},\mu_{i})$ is elementary.
Finally, for $t$ up to $T$, the same reasoning applies after changing the initial condition and time to $t_{0}=t-N^{-1+\omega_{t}}$ and running the dynamics on $[t_{0},t]$ (the local law (4.2) is satisfied at time $t_{0}$ with very high probability, see Proposition 4.13). ∎
4.3 Proof of Proposition 4.5
@@
The following lemma gives some estimates of $\{x_{i}(t,\alpha)\}$, which will be used repeatedly in the rest of this section. We defer its proof in Section 4.4, after the statement of the local law, Proposition 4.13.
@@
Adopt the assumptions in Proposition 4.5, and fix a time ${\mathsf{T}}=\varphi^{-2}$. With very high probability there exists a small constant [^5] In this section the small constant $c_{G}$ may vary from line to line but only depends on $G$ from the $(\nu,G)$-regularity assumption. $c_{G}>0$ such that uniformly for $0\leqslant t\leqslant{\mathsf{T}}$, we have
$c_{G}\,\frac{|i-j|}{N}\leqslant|x_{i}(t,\alpha)-x_{j}(t,\alpha)|\leqslant\frac{|i-j|}{c_{G}\,N},\quad|i-j|\geqslant\varphi^{10},\quad|i|,|j|\leqslant c_{G}N,$ (4.14)
$|x_{i}(t,\alpha)-x_{i}(0,\alpha)|\leqslant\varphi^{3}\left(t+\frac{1}{N}\right),\quad|i|\leqslant c_{G}\,N.$ (4.15)
@@
We need to further introduce the long range and short range operators. Take $c<c_{G}/2$. Denoting $I_{{\mathcal{L}}}=\{(i,j):|i-j|\geqslant\ell,\min\{|i|,|j|\}\leqslant cN\}$, and $I_{{\mathcal{S}}}=\{(i,j):i\neq j\}\setminus I_{{\mathcal{L}}}$, let
@@
$f^{\prime}(t)=-\frac{q}{2}\sum_{(i,j)\in I_{\mathcal{S}}}\mathbb{E}[c_{i,j}(w_{i}-w_{j})(w_{i}^{q-1}-w_{j}^{q-1})]+\sum_{|i|\leqslant N}\frac{q(q-1)}{2}\mathbb{E}\big{[}w_{i}(t)^{q-2}\frac{{\rm d}\langle\xi_{i}\rangle_{t}}{{\rm d}t}\big{]}.$
@@
$\mathbb{E}\left[\sum_{|i|\leqslant N}w_{i}(t)^{q-2}\right]\leqslant 2N\mathbb{E}\left[\Big{(}\frac{1}{2N}\sum_{|i|\leqslant N}w_{i}(t)^{q}\Big{)}^{\frac{q-2}{q}}\right]\leqslant 2N\mathbb{E}\left[\frac{1}{2N}\sum_{|i|\leqslant N}w_{i}(t)^{q}\right]^{\frac{q-2}{q}}=(2N)^{\frac{2}{q}}f(t)^{\frac{q-2}{q}}.$
We have obtained $f^{\prime}(t)\leqslant\frac{p(2p-1)(2N)^{1/p}}{N}f(t)^{\frac{p-1}{p}}$ so with Grönwall’s inequality
@@
Denote the propagator for (4.17) ${\mathcal{U}}={\mathcal{U}}(s,t,\alpha)$ ($s\leqslant t$), i.e. $v_{i}(t)=\sum_{j}{\mathcal{U}}_{ij}(s,t,\alpha)v_{j}(s).$ The propagator preserves the sign (${\mathcal{U}}_{ij}\geqslant 0$) and the mass ($\sum_{j}{\mathcal{U}}_{ij}=1$). The difference of (4.16) and (4.20) gives ${\rm d}(\bm{u}-\bm{w})={\mathcal{B}}(\bm{u}-\bm{w})+{\mathcal{L}}\bm{w}$ so that, with Duhamel’s formula,
@@
For any $|i|\leqslant N$, thanks to Lemma 4.10, we conclude
$\begin{split}\sum_{j}{\mathcal{U}}_{ij}(s,t)({\mathcal{L}}(s)\bm{w}(s))_{j}&\leqslant\sum_{j}{\mathcal{U}}_{ij}(s,t)\sum_{|j-k|\geqslant\ell}\frac{CN}{|j-k|^{2}}|w_{k}(s)|\leqslant\sum_{j}{\mathcal{U}}_{ij}(s,t)\sum_{|j-k|\geqslant\ell}\frac{CN}{|j-k|^{2}}\|\bm{w}(s)\|_{\infty}\\ &\leqslant\sum_{j}{\mathcal{U}}_{ij}(s,t)\frac{CN}{\ell}\|\bm{w}(s)\|_{\infty}\leqslant\frac{CN}{\ell}\|\bm{w}(s)\|_{\infty}\leqslant\frac{CN(\log N)^{2}\sqrt{Ns}}{\ell},\end{split}$ (4.22)
@@
Let ${\bm{w}}$ be the solution of (4.20). Let $\ell=N^{\omega_{\ell}}$ and $\alpha\in[0,1]$. With very high probability, for any $|i|\leqslant\ell$ and $\nu,\omega\ll\omega_{t}\ll\omega_{\ell}<\omega_{K}\ll\tilde{\omega}$, we have
@@
Denote $\zeta=N/\ell$ and $\chi(x)$ be nonnegative, smooth, compactly supported, and such that $\int\chi=1$. In the rest of this proof, we denote, $b_{k}(t)=\alpha b_{k}^{s}(t)+(1-\alpha)b_{k}^{r}(t)$ (from (4.13)). Define $\psi(x)=\int\min\{|x-y|,\frac{c}{2c_{G}}\}\zeta\chi(\zeta y){\rm d}y$ (where $c_{G}$ is introduced in (4.14) and $c$ appears in the definition of $I_{{\mathcal{L}}}$ before (4.18)) and
@@
From the construction $\psi(x)=\frac{c}{2c_{G}}$ for $|x|\geqslant\frac{c}{c_{G}}$. The Itô formula gives
@@
Thus if we define $X_{s}=\sum_{k=1}^{N}v_{k}(s)^{2}$, we obtain
${\rm d}X_{s}=$ $\sum_{k}-2\zeta\psi^{\prime}(x_{k})v_{k}^{2}\frac{{\rm d}b_{k}(s)}{\sqrt{2N}}+2\sum_{k}v_{k}\phi_{k}\frac{{\rm d}\xi_{k}(s)}{\sqrt{2N}}$ (4.24)

### v2

@@ 4 Independence of singular values
For fixed parameters $0<\omega_{K}\ll\widetilde{\omega}$, $K=N^{\omega_{K}}$ we have
$\left|\frac{{\rm d}}{{\rm d}t}\langle b_{i}^{s}-b_{i}^{r},b_{j}^{s}-b_{j}^{r}\rangle_{t}\right|\leqslant N^{-\widetilde{\omega}}\mathds{1}_{1\leqslant|i|,|j|\leqslant K}+4(1-\mathds{1}_{1\leqslant|i|,|j|\leqslant K}).$
We denote the empirical particle density of the initial data $\bm{s}$ as $\mu_{0}=\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\delta_{s_{i}(0)},$ and it Stieltjes transform by $m_{0}(z)$ [^4] Within this section we use $m(z)$ to denote the Stieljes transforms of a measure $\mu$. Instead, we point out that in Theorem 4.1 we used $m^{z}$ to denote the Stieltjes transform of limiting distribution of the singular values of XzX-z.. We assume that $\mu_{0}$ is regular in the following sense.
Assumption 4.3 ($(\nu,R)$-regularity).
Fix $\nu\in(0,1],R>0$, we we recall $\varphi=N^{\nu}$ from (4.1). Let $\widetilde{\mu}$ be a probability measure with Stieltjes transform $\widetilde{m}$ satisfying, for some fixed $C>0$,
$|\widetilde{m}(z)|,|\partial_{z}\widetilde{m}(z)|\leqslant C,\quad C^{-1}\leqslant{\rm Im}[\widetilde{m}(z)]\leqslant C\ \text{ for all }\ |z|\leqslant R.$ (4.8)
We say that $\mu$ is $(\nu,R)$-regular if, for such a $\tilde{\mu}$, its Stieltjes transform $m$ satisfies
$|m(z)-\widetilde{m}(z)|\leqslant\frac{\varphi}{\sqrt{N{\rm Im}[z]}}\ \ \mbox{for any $|z|\leqslant R$ with ${\rm Im}[z]\geqslant\frac{\varphi}{N}$}.$ (4.9)
@@
For our later use, we may deal with the case that $|\widetilde{m}_{0}(z)|,|\partial_{z}\widetilde{m}_{0}(z)|$ may grow with $N$ or ${\rm Im}[\widetilde{m}_{0}(z)]$ may vanish slowly with $N$, namely
$|\widetilde{m}_{0}(z)|,|\partial_{z}\widetilde{m}_{0}(z)|\leqslant N^{\operatorname{o}(1)},{\rm Im}[\widetilde{m}_{0}(z)]\geqslant N^{-\operatorname{o}(1)}.$
@@
Let the processes $\bm{s}(t),\bm{r}(t)$ be solutions of the stochastic differential equations (4.5) and (4.6). Assume that the driving continuous martingales satisfy Assumption 4.2. Additionally assume that the initial data $\bm{s}(0)=\bm{r}(0)$ satisfies $(\nu,R)$-regularity in the sense of Assumption 4.3.
Then, for any choice of small constants $\nu\ll\omega_{t}\ll\omega_{K}\ll\widetilde{\omega}$ there exists $\omega>0$ such that for any $0\leqslant t\leqslant N^{-1+\omega_{t}}$, with very high probability
$|s_{i}(t)-r_{i}(t)|\leqslant N^{-1-\omega},\ \mbox{for all $1\leqslant|i|\leqslant N^{\omega}$}.$
@@
Let the processes $\bm{s}(t),\bm{s}^{\prime}(t)$ both satisfy the stochastic differential equation (4.5), with different initial data and the same driving processes $b_{i}$, which are arbitrary continuous martingales (satisfying (4.7)) and $\bm{s}(0),\bm{s}^{\prime}(0)$ are $(\nu,R)$-regular around ${\rm d}\widetilde{\mu}_{0}(x)=\widetilde{\rho}_{0}(x){\rm d}x,{\rm d}\widetilde{\mu}^{\prime}_{0}(x)=\widetilde{\rho}_{0}^{\prime}(x){\rm d}x$ in the sense of Assumption 4.3, and the two densities are the same at 0: $\widetilde{\rho}_{0}(0)=\widetilde{\rho}^{\prime}_{0}(0)=A$. Then for any $\varepsilon>0$ and $t\geqslant N^{-1+\omega_{t}}$ with $\omega_{t}\gg\nu$, for any $1\leqslant|i|\leqslant N$, with very high probability one has
$|s_{i}(t)-s^{\prime}_{i}(t)|\leqslant N^{\varepsilon}\cdot\frac{|i|}{N}\cdot\left(\frac{1}{\sqrt{Nt}}+\max(\frac{|i|}{N},t)\right).$ (4.10)
@@
Remark 4.8.
In Proposition 4.6, we assumed that the two densities are the same at 0: $\widetilde{\rho}_{0}(0)=\widetilde{\rho}^{\prime}_{0}(0)=A$. If $\widetilde{\rho}_{0}(0)\neq\widetilde{\rho}^{\prime}_{0}(0)=A$, let $a=\widetilde{\rho}_{0}^{\prime}(0)/\widetilde{\rho}_{0}(0)$. We can first do a rescaling for the process $\widehat{s}_{i}(t)=s_{i}(a^{2}t)/a$, the empirical density of $\{\widehat{s}_{i}(0)\}_{1\leqslant|i|\leqslant N}$ is regular around the rescaled density $a\widetilde{\rho}_{0}(ax)$, and $a\widetilde{\rho}_{0}(0)=A=\widetilde{\rho}^{\prime}_{0}(0)$. Moreover, the process $\{\widehat{s}_{i}(t)\}_{1\leqslant|i|\leqslant N}$ satisfies the following system of SDEs:
${\rm d}\widehat{s}_{i}(t)$ $=\frac{1}{\sqrt{2N}}{\rm d}\widehat{b}_{i}^{s}(t)+\frac{1}{2N}\sum_{j\neq i}\frac{1}{\widehat{s}_{i}(t)-\widehat{s}_{j}(t)}{\rm d}t,\qquad\widehat{b}_{i}^{s}(t)=\frac{b_{i}^{s}(a^{2}t)}{a}.$ (4.11)
From the construction, $\langle\widehat{b}^{s}_{i}\rangle_{t}=\langle b^{s}_{i}\rangle_{a^{2}t}/a^{2}$. Thus if ${\rm d}\langle b^{s}_{i}\rangle_{t}/{\rm d}t\leqslant 1$ a.e. as in (4.7), then we also have ${\rm d}\langle\widehat{b}^{s}_{i}\rangle_{t}/{\rm d}t\leqslant 1$ a.e. If Assumption 4.2 holds for the pair $(\{\widehat{b}_{i}^{s}\}_{1\leqslant|i|\leqslant N},\{b_{i}^{r}\}_{1\leqslant|i|\leqslant N})$, then Proposition 4.6 holds for the pair $(\{\widehat{s}_{i}(t)\}_{1\leqslant|i|\leqslant N},\{s^{\prime}_{i}(t)\}_{1\leqslant|i|\leqslant N})$.
Remark 4.9.
The error term $\max(\frac{|i|}{N},t)$ is due to the possibility of $\tilde{\mu}_{0}\neq\tilde{\mu}_{0}^{\prime}$, i.e. it is only due to the propagation of error due to the initial large gaps $s_{j}(0)-s_{j}^{\prime}(0)$, for large $j$. For application such as the Hermitianized matrices considered in this article, $\tilde{\mu}_{0}=\tilde{\mu}_{0}^{\prime}$ and the relaxation statement is simply
$|s_{i}(t)-s^{\prime}_{i}(t)|\leqslant N^{\varepsilon}\cdot\frac{|i|}{N\sqrt{Nt}}.$
@@
We first prove that (4.4) holds for any $t\in[N^{-1+\omega_{t}},2N^{-1+\omega_{t}}]$ for some $\omega,\omega_{t}>0$, and then at the end of the proof we will consider the remaining regime $t\in(2N^{-1+\omega_{t}},\mathcal{T}]$.
@@
${\mathrm{d}}X_{s}=\frac{{\mathrm{d}}B_{s}}{\sqrt{N}},\qquad X_{0}=X.$ (4.12)
Here $B_{s}$ is a matrix valued Brownian motion with entries being i.i.d. standard complex Brownian motions. Fix $z_{l}$, for $l=1,2$, and denote the Hermitization of $X_{s}-z_{l}$ by $H_{s}^{z_{l}}$ (see the analogous definition (3.3)). Then (4.12) induces the following flow on the eigenvalues $x_{i}^{z_{l}}(s)$ of $H_{s}^{z_{l}}$:
@@
with $b_{-i}^{z_{l}}(s)=-b_{i}^{z_{l}}(s)$. Let ${\bm{w}}_{i}^{z_{l}}(s)$ be the eigenvectors of $H_{s}^{z_{l}}$ (they have the form (3.4)), then, for $1\leqslant i,j\leqslant N$, we have
@@
Additionally, by (4.3), it follows that
$\frac{{\rm d}}{{\rm d}s}\langle b_{i}^{z_{1}},b_{j}^{z_{2}}\rangle_{s}\leqslant N^{-2\widetilde{\omega}},\qquad\quad 1\leqslant i,j\leqslant N^{\omega_{K}},$
with very high probability.
By our assumption (4.2) we readily see that the initial condition of (4.12) satisfies Assumption 4.3. It follows that the measure $\zeta$ has a density $\widetilde{\rho}_{0}$. We denote its value at origin by $A=\widetilde{\rho}_{0}(0)$ and we denote the semi-circle distribution by $\rho_{\rm sc}$. Let
$a=\rho_{\rm sc}(0)/A,$ (4.14)
then $\widetilde{\rho}_{0}(x)$ and $\rho_{\rm sc}(x/a)/a$ matches at origin, both equal to $A$.
Denote $\mathbf{b}:=(b_{1}^{z_{1}},\dots,b_{N^{\omega_{K}}}^{z_{1}},b_{1}^{z_{2}},\dots,b_{N^{\omega_{K}}}^{z_{2}})\in\mathbb{R}^{2N^{\omega_{K}}}$, for simplicity of notation we assume that $N^{\omega_{K}}$ is an integer. Next, by the martingale representation theorem (see e.g. [Kal02, Theorem 18.12]) we have
${\mathrm{d}}\mathbf{b}(t)=\sqrt{C(t)}{\mathrm{d}}{\bm{\beta}}(t),$
where $C(t)$ is the $(2N^{\omega_{K}})\times(2N^{\omega_{K}})$ covariance matrix of the vector $\mathbf{b}(t)$, and ${\bm{\beta}}(t)\in\mathbb{R}^{2N^{\omega_{K}}}$ is a vector of standard real i.i.d. Brownian motions. We now consider two fully independent processes $y_{i}^{(l)}(t)$, for $l=1,2$, with $y_{i}^{(l)}(t)$ the eigenvalues of the Hermitization of a Ginibre matrix rescaled by $a$ (recall from (4.14)), satisfying (4.13) with $b_{i}^{z_{l}}(t)$ being replaced by $\beta_{i}^{(l)}$, where $\beta_{i}^{(1)}$, $\beta_{i}^{(2)}$ are a family of $2N$ i.i.d. standard real Brownian motions such that $\beta_{i}^{(1)}=({\bm{\beta}})_{i}$, for $1\leqslant i\leqslant N^{\omega_{K}}$, and $\beta_{i-N^{\omega_{K}}}^{(2)}=({\bm{\beta}})_{i}$, for $N^{\omega_{K}}+1\leqslant i\leqslant 2N^{\omega_{K}}$. For these martingales, with very high probability, we thus have
$\left|\frac{{\mathrm{d}}}{{\mathrm{d}}t}\langle\mathbf{b}_{i}-{\bm{\beta}}_{i},\mathbf{b}_{j}-{\bm{\beta}}_{j}\rangle_{t}\right|=\left[\left(\sqrt{C(t)}-I\right)^{2}\right]_{ij}\leqslant\frac{N^{2\omega_{K}}}{N^{4\tilde{\omega}}}\leqslant N^{-\tilde{\omega}}$,
where in the last inequality we used that $\omega_{K}\leqslant\tilde{\omega}/10$ (see the display below (4.1)). This shows that $\{b_{i}^{z_{l}}\}_{i}$, $\{\beta_{i}^{(l)}\}_{i}$ satisfy Assumption 4.2. Additionally, by our assumption (4.2) we readily see that the initial condition of (4.12) satisfies Assumption 4.3. Moreover, the initial condition of $y_{i}^{(l)}(t)$ also satisfies Assumption 4.3 with respect to the scaled semi-circle distribution $\rho_{\rm sc}(x/a)/a$. And by (4.14) $\widetilde{\rho}(0)=A=\rho_{\rm sc}(0)/a$.
We can thus apply Propositions 4.54.6 to see that
$\big|x_{i}^{z_{l}}(t)-y_{i}^{(l)}(t)\big|\leqslant N^{-1-\omega},\qquad\quad 1\leqslant|i|\leqslant N^{\omega},$ (4.15)
with very high probability. More precisely, to obtain (4.15), we first used Proposition 4.5 to show that $|x_{i}^{z_{l}}(t)-\widetilde{y}_{i}^{(l)}(t)|\leqslant N^{-1-\omega}$, where $\widetilde{y}_{i}^{(l)}(t)$ is the solution of
${\mathrm{d}}\widetilde{y}_{i}^{(l)}(s)=\frac{{\mathrm{d}}\beta_{i}^{(l)}(t)}{\sqrt{2N}}+\frac{1}{2N}\sum_{j\neq i}\frac{1}{\widetilde{y}_{i}^{(l)}(s)-\widetilde{y}_{j}^{(l)}(s)}{\mathrm{d}}s$
with initial data $\widetilde{y}_{i}^{(l)}(0)=x_{i}^{z_{l}}(0)$, and for $\beta_{i}^{(l)}$ as above. This first step does not change the initial data but replaces the driving martingales in (4.13) with i.i.d. standard real Brownian motions. Then, in the second step, we used Proposition 4.6 to show $|\widetilde{y}_{i}^{(l)}(t)-y_{i}^{(l)}(t)|\leqslant N^{-1-\omega}$. These two steps together give (4.15). We notice that $\{y_{i}^{(l)}(t)\}_{1\leqslant|i|\leqslant N}$ have the same law as the eigenvalues of the Hermitization of a Ginibre matrix rescaled by $\sqrt{a^{2}+t^{2}}=(1+\operatorname{o}(1))\rho_{\rm sc}(0)/\widetilde{\rho}_{0}(0)$. The transfer from (4.15) on the (signed) eigenvalues $(x_{i},y_{i})$ to the singular values $(\lambda_{i},\mu_{i})$ is elementary. This leads to the claim (4.4) for $t\in[N^{-1+\omega_{t}},2N^{-1+\omega_{t}}]$.
Finally, for $t$ up to $\mathcal{T}$, the same reasoning applies after changing the initial condition and time to $t_{0}=t-N^{-1+\omega_{t}}$ and running the dynamics on $t_{0},t]$ (the local law (4.2) is satisfied at time $t_{0}$ with very high probability, see Proposition 4.15). ∎
4.3 Proof of Proposition 4.5.
@@
The following lemma gives some estimates of $\{x_{i}(t,\alpha)\}$, which will be used repeatedly in the rest of this section. We defer its proof to Section 4.4, after the statement of the local law, Proposition 4.15.
@@
Adopt the assumptions in Proposition 4.5, and fix a time ${\mathsf{T}}=\varphi^{-2}$. With very high probability there exists a small constant [^5] In this section the small constant $c_{R}$ may vary from line to line but only depends on $R$ from the $(\nu,R)$-regularity assumption. $c_{R}>0$ such that uniformly for $0\leqslant t\leqslant{\mathsf{T}}$, we have
$c_{R}\,\frac{|i-j|}{N}\leqslant|x_{i}(t,\alpha)-x_{j}(t,\alpha)|\leqslant\frac{|i-j|}{c_{R}\,N},\quad|i-j|\geqslant\varphi^{10},\quad|i|,|j|\leqslant c_{R}N,$ (4.17)
$|x_{i}(t,\alpha)-x_{i}(0,\alpha)|\leqslant\varphi^{3}\left(t+\frac{1}{N}\right),\quad|i|\leqslant c_{R}\,N.$ (4.18)
@@
We need to further introduce the long range and short range operators. Take $c<c_{R}/2$. Denoting $I_{{\mathcal{L}}}=\{(i,j):|i-j|\geqslant\ell,\min\{|i|,|j|\}\leqslant cN\}$, and $I_{{\mathcal{S}}}=\{(i,j):i\neq j\}\setminus I_{{\mathcal{L}}}$, let
@@
$\partial_{t}f(t)=-\frac{q}{2}\sum_{(i,j)\in I_{\mathcal{S}}}\mathbb{E}[c_{i,j}(w_{i}-w_{j})(w_{i}^{q-1}-w_{j}^{q-1})]+\sum_{1\leqslant|i|\leqslant N}\frac{q(q-1)}{2}\mathbb{E}\big[w_{i}(t)^{q-2}\frac{{\rm d}\langle\xi_{i}\rangle_{t}}{{\rm d}t}\big].$
@@
$\mathbb{E}\left[\sum_{1\leqslant|i|\leqslant N}w_{i}(t)^{q-2}\right]\leqslant 2N\mathbb{E}\left[\Big(\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}w_{i}(t)^{q}\Big)^{\frac{q-2}{q}}\right]\leqslant 2N\mathbb{E}\left[\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}w_{i}(t)^{q}\right]^{\frac{q-2}{q}}=(2N)^{\frac{2}{q}}f(t)^{\frac{q-2}{q}}.$
We have obtained $\partial_{t}f(t)\leqslant\frac{p(2p-1)(2N)^{1/p}}{N}f(t)^{\frac{p-1}{p}}$ so with Grönwall’s inequality
@@
Denote the propagator for (4.20) ${\mathcal{U}}={\mathcal{U}}(s,t,\alpha)$ ($s\leqslant t$), i.e. $v_{i}(t)=\sum_{1\leqslant|j|\leqslant N}{\mathcal{U}}_{ij}(s,t,\alpha)v_{j}(s).$ The propagator preserves the sign (${\mathcal{U}}_{ij}\geqslant 0$) and the mass ($\sum_{1\leqslant|j|\leqslant N}{\mathcal{U}}_{ij}=1$). The difference of (4.19) and (4.23) gives ${\rm d}(\bm{u}-\bm{w})={\mathcal{B}}(\bm{u}-\bm{w})+{\mathcal{L}}\bm{w}$ so that, with Duhamel’s formula,
@@
For any $1\leqslant|i|\leqslant N$, thanks to Lemma 4.12, we conclude
$\begin{split}\sum_{1\leqslant|j|\leqslant N}{\mathcal{U}}_{ij}(s,t)({\mathcal{L}}(s)\bm{w}(s))_{j}&\leqslant\sum_{1\leqslant|j|\leqslant N}{\mathcal{U}}_{ij}(s,t)\sum_{|j-k|\geqslant\ell}\frac{CN}{|j-k|^{2}}|w_{k}(s)|\leqslant\sum_{1\leqslant|j|\leqslant N}{\mathcal{U}}_{ij}(s,t)\sum_{|j-k|\geqslant\ell}\frac{CN}{|j-k|^{2}}\|\bm{w}(s)\|_{\infty}\\ &\leqslant\sum_{1\leqslant|j|\leqslant N}{\mathcal{U}}_{ij}(s,t)\frac{CN}{\ell}\|\bm{w}(s)\|_{\infty}\leqslant\frac{CN}{\ell}\|\bm{w}(s)\|_{\infty}\leqslant\frac{CN(\log N)^{2}\sqrt{Ns}}{\ell},\end{split}$ (4.25)
@@
Let $\bm{w}$} be the solution of (4.23). Let $\ell=N^{\omega_{\ell}}$ and $\alpha\in[0,1]$. With very high probability, for any $1\leqslant|i|\leqslant\ell$, $\nu,\omega\ll\omega_{t}\ll\omega_{\ell}<\omega_{K}\ll\widetilde{\omega}$ and $t=N^{-1+\omega_{t}}$, we have
@@
Denote $\zeta=N/\ell$ and $\chi(x)$ be nonnegative, smooth, compactly supported, and such that $\int\chi=1$. In the rest of this proof, we denote, $b_{k}(t)=\alpha b_{k}^{s}(t)+(1-\alpha)b_{k}^{r}(t)$ (from (4.16)). Define $\psi(x)=\int\min\{|x-y|,\frac{c}{2c_{R}}\}\zeta\chi(\zeta y){\rm d}y$ (where $c_{R}$ is introduced in (4.17) and $c$ appears in the definition of $I_{{\mathcal{L}}}$ before (4.21)) and
@@
From the construction $\psi(x)=\frac{c}{2c_{R}}$ for $|x|\geqslant\frac{c}{c_{R}}$. The Itô formula gives
@@
Thus if we define $X_{s}=\sum_{1\leqslant|k|\leqslant N}v_{k}(s)^{2}$, we obtain
${\rm d}X_{s}=$ $\sum_{1\leqslant|k|\leqslant N}-2\zeta\psi^{\prime}(x_{k})v_{k}^{2}\frac{{\rm d}b_{k}(s)}{\sqrt{2N}}+2\sum_{1\leqslant|k|\leqslant N}v_{k}\phi_{k}\frac{{\rm d}\xi_{k}(s)}{\sqrt{2N}}$ (4.27)